# Lifted Forward Planning in Relational Factored Markov Decision Processes with Concurrent Actions

### v2

@@ Abstract.
When allowing concurrent actions in Markov Decision Processes, whose state and action spaces grow exponentially in the number of objects, computing a policy becomes highly inefficient, as it requires enumerating the joint of the two spaces. For the case of indistinguishable objects, we present a first-order representation to tackle the exponential blow-up in the action and state spaces. We propose Foreplan, an efficient relational forward planner, which uses the first-order representation allowing to compute policies in space and time polynomially in the number of objects. Thus, Foreplan significantly increases the number of planning problems solvable, in an exact manner, in reasonable time, which we underscore with a theoretical analysis. To speed up the computations even further, we also introduce an approximate version of Foreplan, including guarantees on the error. Further, we provide an empirical evaluation of both Foreplan versions, demonstrating a speedup of several orders of magnitude. For the approximate version of Foreplan, we also empirically show that the induced error is often negligible.
@@
Planning, Markov Decision Process, Lifting, Probabilistic Graphical Models
@@ 1. Introduction
To compute a policy for a Markov decision processes (MDPs), combinations of the state and action spaces have to be enumerated to find the best possible action for each state. With increasing number of (indistinguishable) objects, the state space grows exponentially, leading to drastically more combinations to be enumerated. In case the actions also depend on the states or objects, the action space grows as well, leading to even more combinations for enumeration. Further, if one allows concurrent actions, all possible combinations of the actions, yielding an exponentially-sized action space, of the already enlarged action space have to be used during the enumeration. Therefore, given the number of possibilities that have to be enumerated, concurrency is rarely modelled. However, consider the following example: A small town is haunted by an epidemic. To fight the epidemic, the town’s mayor can impose travel bans on the town’s citizens. Certainly, the mayor can confine all citizens to their homes, stopping the epidemic. However, the citizens’ overall welfare is important as well. Therefore, the mayor is interested in the best decision of imposing travel bans (concurrent actions) w.r.t. confining the epidemic, while keeping the citizens’ welfare above a certain threshold. In addition, there can be circumstances where the policy has to be adapted, e.g., if a wave of infections is on the rise, threatening to infect the majority of the population. Computing this problem on a propositional level, with every citizen explicitly represented, like approaches such as MDPs do, blows up the state and action space, as it requires exponential enumeration of all subsets of the population. However, there are groups of citizens behaving identically w.r.t. getting sick (and well again) as well as regarding their welfare if a travel ban is imposed, i.e., these citizens are indistinguishable for the mayor. Within these groups, it does not matter on which exact citizen the mayor imposes a travel ban on, only on how many she imposes a travel ban on. Additionally, many computations over subsets of indistinguishable citizens are redundant. Thus, we propose to drastically reduce the search and action space by using a first-order representation, grouping indistinguishable citizens, and use this representation to plan group-wise by reasoning about a representative and then projecting the result to the whole group.
@@
First, to be able to represent numerous indistinguishable objects and actions, we use probabilistic relational models as a representation in (factored) MDPs (fMDPs), which yields relational factored MDPs (rfMDPs). Second, as our key contribution, we propose Foreplan to carry out exact planning in rfMDPs. Foreplan uses our first-order representation to keep the action and state space only polynomial in the number of objects, which we prove by using our so-called relational cost graph. Foreplan remains exponential in the number of cliques $c$ and size $w$ of the largest clique in the relational cost graph, but both parameters are mostly small and independent of the domain sizes. Therefore, Foreplan is an efficient planner w.r.t the number of objects. We show a speedup of at least four orders of magnitude is achievable. Last, following the approximation ideas of approx_factored_mdps, we can reduce the runtime even further. We propose Approximate Foreplan, whose runtime is polynomial in $c$. For Approximate Foreplan, we show a speedup of at least four orders of magnitude.
@@
mdp introduces MDPs, which factored_mdp extend to fMDPs by factorizing the transition function. Factorizing also the value function, approx_factored_mdps provide two approximate algorithms for solving planning in fMDPs. dean2013model cluster the state space of fMDPs to reduce the state space even further. givan2003equivalence group equivalent states based on the notion of bisimulation. Both approaches lack the ability to handle concurrent actions efficiently. MDPs can be generalized to partially observable MDPs, in which the agent is uncertain in which state the environment currently is kaelbling1998planning. sanner2010symbolic add the first-order perspective to partially observable MDPs, focusing on the observations and without concurrent actions. bernstein2002complexity extend partially observable MDPs to have a set of decentralized agents. braun2022lifting work with groups of indistinguishable agents. This is similar to our approach, in which we handle sets of indistinguishable state variables. However, they do not provide an algorithm.
The idea of lifting is to carry out computations over representatives for groups of indistinguishable random variables lve; kersting2012lifted; raedt2016statistical. There are online decision making approaches adding action and utilitity nodes to this representation apsel2011extended; 10.1007/978-3-030-12738-1_10; 10.1007/978-3-030-18305-9_33, here, we focus on offline planning. To carry out even more lifted computations, crv extends lifted probabilistic inference by a generalized counting framework, which we extend later on. Using a first-order representation for states, actions and objects, boutilier2001symbolic exploit the relational structure for planning using MDPs, still without concurrent actions. To specify factored transition models in first-order MDPs, sanner2007approximate introduce factored first-order MDPs, which uses a backward search and provide only an approximate algorithm. In contrast, we propose an exact algorithm by applying lifting. Later, we also introduce an approximate version. Moreover, we prove on which models our algorithm runs efficiently. For a survey on planning using first-order representations, we refer to correa2024lifted.
@@ 2. Preliminaries
In this section, we lay the foundation for rfMDPs. We first recap (f)MDPs, which model probabilistic state changes, when an agent performs an action. Furthermore, the states have a reward assigned, and the task is to compute the optimal policy, i.e., which action to perform in which state, for the agent w.r.t. rewards. Second, we recap lifting in probabilistic graphical models.
@@
A Markov Decision Process is a tuple $(\mathbf{S},\mathbf{A},T,R)$ with a set of states $\mathbf{S}$, a set of actions $\mathbf{A}$, a reward function $R:\mathbf{S}\mapsto\mathbb{R}$ and a transition function $T:\mathbf{S}\times\mathbf{A}\times\mathbf{S}\to[0,1]$. The value $T(s,a,s^{\prime})$ is the probability $P(s^{\prime}\mid s,a)$ of transitioning from state $s\in\mathbf{S}$ to state $s^{\prime}\in\mathbf{S}$ if action $a\in\mathbf{A}$ is performed.
The rewards are additive, possibly discounted by a factor $\gamma\in[0,1)$. An MDP is fully observable and has the first order Markov property, i.e., the probability of the next state only depends on the current state and action. Let us have a look at a simple, (incomplete) example of an MDP.
@@
Planning in MDPs refers to calculating an optimal policy for the agent, which is a mapping from each state to an action to perform. To compute such a policy, we first define the utility of a state:
@@
To find the utility of a state algorithmically, we find a value function $V$ satisfying the Bellman equation. The value function induces a policy by selecting the action that yields the maximum expected value. For computing a value function, we can use a linear programming formulation mdp_lp_orig; mdp_lp:
@@
where the coefficients $\alpha(s)$ are arbitrary positive numbers, e.g., an equal distribution over all states mdp_lp. Planning in MDPs can be solved in polynomial time w.r.t to the state space size papadimitriou1987complexity. But what if the state space becomes very large, e.g., exponential in the number of objects? For retaining an efficient transition model, fMDPs make use of state variables for the objects. The state space is then spanned by the state variables and is exponential in the number of state variables. For simplicity, all state variables are Boolean, but can be easily extended to non-Boolean.
@@
When having numerous objects, the state space explodes. But when these objects are indistinguishable, we can use parameterized graphical models to encode the state space and transition model compactly.
@@ 2.2. Parameterized Graphical Models
Having indistinguishable random variables leads to the redundant computations. We can tackle redundant computations by parameterizing our probabilistic model and grouping indistinguishable variables, so that inference in the probabilistic model becomes tractable w.r.t. domain sizes by using representatives during calculations niepert2014tractability. We first define parameterized random variables (PRVs) to group same behaving variables:
@@
We create a PRV for the health status of each person in the town:
@@
Let $\mathbf{W}=\{Sick,Epidemic\}$, $\mathbf{L}=\{M\}$, $\mathcal{D}(M)=\mathbf{D}=\{a,b,c,d,e,f,g,h\}$ with Boolean-valued PRVs $Sick(M)$ and $Epidemic$.
EpidemicEpidemic is a parameterless PRV. To set PRVs into relation, we use parameterized factors:
@@
Let $\Phi$ be a set of factor names. We denote a parameterized factor (parfactor) $g$ by $\phi(\mathcal{B})_{|C}$ with $\mathcal{B}=(B_{1},\dots,B_{n})$ a sequence of PRVs, $\phi:\times_{i=1}^{n}\mathcal{R}(B_{i})\to\mathbb{R}^{+}$ a potential function with name $\phi\in\Phi$, and $C$ a constraint on the logvars of $\mathcal{B}$. A set of parfactors forms a model $G:=\{g_{i}\}_{i=1}^{n}$. With $Z$ as normalizing constant, $G$ represents the full joint distribution $P_{G}=\frac{1}{Z}\prod_{f\in gr(G)}f$, with $gr(G)$ referring to the groundings of $G$ w.r.t. given constraints. A grounding is the instantiation of each parfactor with an allowed constant.
@@
Let $\forall m\in M:\phi(Sick(m),Epidemic)_{\top}$ be the parfactor with potential $\phi$ defining the probability to be sick for all persons $\mathbf{D}$, given there is an epidemic (or not). The grounded model then consists of the eight factors $\phi(Sick(a),Epidemic)$, $\phi(Sick(b)$, $Epidemic)$, …, $\phi(Sick(h),Epidemic)$ compared to one parfactor in the lifted model. Table 1 shows the potential function.
@@
Next, we integrate the parfactor into fMDPs to achieve a compact representation of the state and action spaces.
@@ 3. Relational Factored MDPs
We present the first-order representation for the state and action space as well as the transition function, which allows to efficiently represent indistinguishable objects as well as concurrent action, which then in turn can be exploited by an appropriate algorithm. To that end, we present Foreplan in the next section.
We define rfMDPs based on fMDPs, but include a parfactor model inside the state and action space and transition function. That is, the state variables can now contain PRVs, which are then used in the transition model and the reward function. Also, we have action PRVs, whose value is chosen by the agent. Basically, we keep the same semantics as in Definition 2.1 and only change the representation to exploit indistinguishable variables. In the following definition we use the term interpretation of a set for a truth-value assignment to each element in the set.
@@
A relational factored MDP is a tuple $(\mathbf{D},\mathbf{L},\mathbf{X},\mathbf{A},G,\mathbf{R})$. The set $\mathbf{D}$ is a set of constants and the set $\mathbf{L}$ is a set of logvars over $\mathbf{D}$. The set $\mathbf{X}$ is a set of PRVs defined over $\mathbf{L}$. The set of possible interpretations $\mathcal{I}_{\mathbf{X}}$ for the groundings of the set $\mathbf{X}$ defines the state space, i.e., all possible assignments. The set $\mathbf{A}$ is a set of action PRVs. A parfactor model $G$ over $\mathbf{A}$ and $\mathbf{X}$ represents the transition function $T:\mathcal{I}_{\mathbf{X}}\times\mathcal{I}_{\mathbf{A}}\times\mathcal{I}_{\mathbf{X}}\to\mathbb{R}_{0}^{+}$, with the set $\mathcal{I}_{\mathbf{A}}$ of possible interpretations of the groundings of $\mathbf{A}$, and specifies the transition probability given an action and a previous state. The set $\mathbf{R}$ contains parameterized local reward functions $R_{i}:\bigtimes_{j}\mathcal{R}(B_{i,j})\to\mathbb{R}$, defined over all PRVs $B_{i,j}$, which contribute to the local reward $R_{i}$. The reward function $R$ is decomposed as a sum over $\mathbf{R}$.
@@
Grounded graphical model for the epidemic example: The factor f_1 links Travel(a) and Restrict(a) to Travel’(a), analogously tilde f_1 for b, f_2 links Sick(a) and Epidemic to Sick’(a), analogously tilde f_2 for b, and f_3 links Travel(a) and Travel(b) to Epidemic’.
@@
Figure 2. Lifted representation of the transition model for Example 3. We abbreviate by using only the first letter(s) for each symbol.
@@
Since an action can be applied to each person concurrently, the amount of possible actions is exponential due to the power set, i.e., all possible combinations for all persons. The action PRVs help Foreplan to reduce the exponential blow-up by exploiting the indistinguishability.
With Definition 3, we require the rewards to be represented as a sum of local reward functions, and not as some arbitrary function. Actually, this is too restrictive: It is sufficient that an operation running over same behaving variables has a hyperoperation repeating the operation multiple times. For a sum, the hyperoperation is multiplication. Also, multiple terms of such operations can be linked together in any way. In Appendix A.1, we give examples for hyperoperations. For simplicity, we use the definition as it is throughout the paper.
While Definition 3 defines the ground semantics, there are symmetries within the parfactor model, which we exploit with Foreplan in the next section. In the remainder of this section, we explain what we mean by action PRVs and by parameterized local reward functions.

### v3

@@ Abstract.
When allowing concurrent actions in Markov Decision Processes, whose state and action spaces grow exponentially in the number of objects, computing a policy becomes highly inefficient, as it requires enumerating the joint of the two spaces. For the case of indistinguishable objects, we present a first-order representation to tackle the exponential blow-up in the action and state spaces. We propose Foreplan, an efficient relational forward planner, which uses the first-order representation allowing to compute policies in space and time polynomially in the number of objects. Thus, Foreplan significantly increases the number of planning problems solvable in an exact manner in reasonable time, which we underscore with a theoretical analysis. To speed up computations even further, we also introduce an approximate version of Foreplan, including guarantees on the error. Further, we provide an empirical evaluation of both Foreplan versions, demonstrating a speedup of several orders of magnitude. For the approximate version of Foreplan, we also empirically show that the induced error is often negligible.
@@
Planning; Markov Decision Process; Lifting; Probabilistic Graphical Models
doi:
GOIS4183
@@ 1. Introduction
To compute a policy for a Markov decision processes (MDPs), combinations of the state and action spaces have to be enumerated to find the best possible action for each state. With an increasing number of (indistinguishable) objects, the state space grows exponentially, leading to drastically more combinations to be enumerated. In case the actions also depend on the states or objects, the action space grows as well, leading to even more combinations to enumerate. Further, if one allows concurrent actions, all possible combinations of the actions of the already enlarged action space have to be accounted for during the enumeration, yielding an exponentially-sized action space. Therefore, given the number of combinations that have to be enumerated, concurrency is rarely modelled. However, consider the following example: A small town is haunted by an epidemic. To fight the epidemic, the town’s mayor can impose travel bans on the town’s citizens. Certainly, the mayor can confine all citizens to their homes, stopping the epidemic. However, the citizens’ overall welfare is important as well. Therefore, the mayor is interested in the best decision of imposing travel bans (concurrent actions) w.r.t. confining the epidemic, while keeping the citizens’ welfare above a certain threshold. Computing this problem on a propositional level, with every citizen explicitly represented, such as with current MDP approaches, blows up the state and action space, as it requires exponential enumeration of all subsets of the population. However, there are groups of citizens behaving identically w.r.t. getting sick (and well again) as well as regarding their welfare if a travel ban is imposed, i.e., these citizens are indistinguishable for the mayor. Within these groups, it does not matter on which exact citizen the mayor imposes a travel ban, only on how many she imposes a travel ban is important. Additionally, many computations over subsets of indistinguishable citizens are redundant. Thus, we propose to drastically reduce the search and action space using a first-order representation, grouping indistinguishable citizens, and use this representation to plan group-wise by reasoning about a representative and then projecting the result to the whole group. This paper is an extended version of our paper accepted at AAMAS 2026 aamas with appendix including extended examples and with all (complete) proofs.
@@
First, to be able to represent numerous indistinguishable objects and actions, we use probabilistic relational models as a representation in (factored) MDPs (fMDPs), which yields relational factored MDPs (rfMDPs). Second, as our key contribution, we propose Foreplan to carry out exact planning in rfMDPs. Foreplan uses our first-order representation to keep the action and state space only polynomial in the number of objects, which we prove by using our so-called relational cost graph. Foreplan remains exponential in the number of cliques $c$ and size $w$ of the largest clique in the relational cost graph, but both parameters are mostly small and independent of the domain sizes. Therefore, Foreplan is an efficient planner w.r.t the number of objects. We show a speedup of at least four orders of magnitude is achievable. Last, following the approximation ideas of approx_factored_mdps, we can reduce the runtime even further. We propose Approximate Foreplan, whose runtime is polynomial in $c$ unlike Foreplan. For Approximate Foreplan, we show a speedup of at least four orders of magnitude.
@@
mdp introduces MDPs, which factored_mdp extend to fMDPs by factorizing the transition function. Factorizing also the value function, approx_factored_mdps provide two approximate algorithms for solving planning in fMDPs. dean2013model cluster the state space of fMDPs to reduce the state space even further. givan2003equivalence group equivalent states based on the notion of bisimulation. Both approaches lack the ability to handle concurrent actions efficiently. MDPs can be generalized to partially observable MDPs, in which the agent is uncertain in which state the environment currently is kaelbling1998planning. sanner2010symbolic add the first-order perspective to partially observable MDPs, but do not consider concurrent actions. bernstein2002complexity extend partially observable MDPs to have a set of decentralized agents. braun2022lifting group indistinguishable agents. This is similar to our approach, in which we handle sets of indistinguishable state variables. However, braun2022lifting only solve the problem inefficiently via brute-force.
The idea of lifting is to carry out computations over representatives for groups of indistinguishable random variables lve; kersting2012lifted; raedt2016statistical. There are online decision making approaches adding action and utility nodes to this representation apsel2011extended; 10.1007/978-3-030-12738-1_10; 10.1007/978-3-030-18305-9_33, here, we focus on offline planning. To carry out even more lifted computations, crv extends lifted probabilistic inference by a generalized counting framework, which we extend later on. Using a first-order representation for states, actions and objects, boutilier2001symbolic exploit the relational structure for planning using MDPs, still without concurrent actions. To specify factored transition models in first-order MDPs, sanner2007approximate introduce factored first-order MDPs using a backward search, and give only an approximate algorithm without providing error bounds. In contrast, we propose an exact algorithm by applying lifting. Later, we also introduce an approximate version with error estimation. Moreover, we prove on which models our algorithm runs efficiently. For a survey on planning using first-order representations, we refer to correa2024lifted.
@@ 2. Preliminaries
In this section, we lay the foundation for rfMDPs. We first recap (f)MDPs, which model probabilistic state changes occurring due to agents performing actions. Furthermore, states have a reward assigned, and the task is to compute the optimal policy, i.e., which action to perform in which state. Second, we recap lifting in probabilistic graphical models.
@@
A Markov Decision Process is a tuple $(\mathbf{S},\mathbf{A},T,R)$ with a finite set of states $\mathbf{S}$, a finite set of actions $\mathbf{A}$, a reward function $R:\mathbf{S}\mapsto\mathbb{R}$ and a transition function $T:\mathbf{S}\times\mathbf{A}\times\mathbf{S}\to[0,1]$. The value $T(s,a,s^{\prime})$ is the probability $P(s^{\prime}\mid s,a)$ of transitioning from state $s\in\mathbf{S}$ to state $s^{\prime}\in\mathbf{S}$ if action $a\in\mathbf{A}$ is performed.
The rewards are additive, possibly discounted by a factor $\gamma\in[0,1)$. An MDP is fully observable and has the first order Markov property, i.e., the probability of the next state only depends on the current state and action. Let us have a look at a simple, yet incomplete example of an MDP.
@@
Planning in MDPs refers to calculating an optimal policy, which is a mapping from each state to an action to perform for the agent. To compute such a policy, we first define the utility of a state:
@@
To find the utility of a state algorithmically, we determine a value function $V$ satisfying the Bellman equation. The value function induces a policy by selecting the action that yields the maximum expected value. For computing a value function, we can use a linear programming formulation mdp_lp_orig; mdp_lp:
@@
where the coefficients $\alpha(s)$ are arbitrary positive numbers, e.g., a uniform distribution over all states mdp_lp. Planning in MDPs can be solved in polynomial time w.r.t to the state space size papadimitriou1987complexity. But what if the state space becomes very large, e.g., exponential in the number of objects? For retaining an efficient transition model, fMDPs make use of state variables for the objects. The state space is then spanned by the state variables. For simplicity, all state variables are Boolean, but can be easily extended to non-Boolean.
@@
For large $m$, the state space explodes. But when these objects are indistinguishable, we can use parameterized graphical models to encode state spaces and transition models in a compact way.
@@ 2.2. Parameterized Graphical Models
Having indistinguishable random variables leads to redundant computations. We can tackle redundant computations by parameterizing our probabilistic model and using representatives of groups of indistinguishable variables, so that inference in the probabilistic model becomes tractable w.r.t. domain sizes using representatives during calculations niepert2014tractability. We first define parameterized random variables (PRVs) to group same behaving variables:
@@
We create PRVs for the health status of each person in the town:
@@
Let $\mathbf{W}=\{Sick,Epidemic\}$, $\mathbf{L}=\{M\}$, $\mathcal{D}(M)=\mathbf{D}=\{a,b,c,d,e,f,g,h\}$ with Boolean-valued PRVs $Sick(M)$ and $Epidemic$. EpidemicEpidemic is a parameterless PRV.
To relate PRVs, we use parameterized factors:
@@
Let $\Phi$ be a set of factor names. We denote a parameterized factor (parfactor) $g$ by $\phi(\mathcal{B})_{|C}$ with $\mathcal{B}=(B_{1},\dots,B_{n})$ a sequence of PRVs, $\phi:\times_{i=1}^{n}\mathcal{R}(B_{i})\to\mathbb{R}^{+}$ a potential function with name $\phi\in\Phi$, and $C$ a constraint on the logvars of $\mathcal{B}$. A set of parfactors forms a model $G:=\{g_{i}\}_{i=1}^{n}$. With $Z$ as normalizing constant, $G$ represents the full joint distribution $P_{G}=\frac{1}{Z}\prod_{f\in gr(G)}f$, with $gr(G)$ referring to the groundings of $G$ w.r.t. given constraints. A grounding is the instantiation of each logvar in each parfactor with an allowed constant.
@@
Let $\forall m\in M:\phi(Sick(m),Epidemic)_{\top}$ be the parfactor with potential $\phi$ defining the probability to be sick for all persons from $\mathbf{D}$, given there is an epidemic (or not). The grounded model then consists of the eight factors $\phi(Sick(a),Epidemic)$, …, $\phi(Sick(h),Epidemic)$ compared to one parfactor in the lifted model. Table 1 shows the potential function.
@@
Next, we integrate parfactors into fMDPs to achieve a compact representation of state and action spaces.
@@ 3. Relational Factored MDPs
Now, we present the first-order representation for the state and the action space as well as the transition function, which allows to efficiently represent indistinguishable objects as well as concurrent action, which then in turn can be exploited by an appropriate algorithm. To that end, we present Foreplan in the next section.
We define rfMDPs based on fMDPs, but include a parfactor model inside the state and action space and transition function. That is, the set of state variables can now contain PRVs, which are then used in the transition model and the reward function. Also, we have action PRVs, whose value is chosen by the agent. Basically, we keep the same semantics as in Definition 2.1 and only change the representation to exploit indistinguishable variables. In the following definition we use the term interpretation of a set for a truth-value assignment to each element in the set.
@@
A relational factored MDP is a tuple $(\mathbf{D},\mathbf{L},\mathbf{B},\mathbf{A},G,\mathbf{R})$. The set $\mathbf{D}$ is a set of constants and the set $\mathbf{L}$ is a set of logvars over $\mathbf{D}$. The set $\mathbf{B}$ is a set of PRVs defined over $\mathbf{L}$. The set of possible interpretations $\mathcal{I}_{\mathbf{B}}$ for the groundings of the set $\mathbf{B}$ defines the state space, i.e., all possible assignments. The set $\mathbf{A}$ is a set of action PRVs. A parfactor model $G$ over $\mathbf{A}$ and $\mathbf{B}$ represents the transition function $T:\mathcal{I}_{\mathbf{B}}\times\mathcal{I}_{\mathbf{A}}\times\mathcal{I}_{\mathbf{B}}\to\mathbb{R}_{0}^{+}$, with the set $\mathcal{I}_{\mathbf{A}}$ of possible interpretations of the groundings of $\mathbf{A}$, and specifies the transition probability given an action and a previous state. The set $\mathbf{R}$ contains parameterized local reward functions $R_{i}:\bigtimes_{j}\mathcal{R}(B_{i,j})\to\mathbb{R}$, defined over all PRVs $B_{i,j}$, which contribute to the local reward $R_{i}$. The reward function $R$ is decomposed as a sum operation over $\mathbf{R}$.
@@
Grounded graphical model for the epidemic example: The factor f_1 links Travel(a) and Reestrict(a) to Travel’(a), analogously ~f_1 for b, f_2 links Sick(a) and Epidemic to Sick’(a), analogously ~f_2 for b, and f_3 links Travel(a) and Travel(b) to Epi’.
@@
Figure 2. Lifted representation of the transition model for Example 3. We abbreviate by using only the first letter(s) for each symbol. Figure 1 shows the lifted representation.
@@
Parameterized graphical model for the epidemic example: The factor f_1 links Travel(M) and Restrict(M) to Travel’(M), f_2 links Sick(M) and Epidemic to Sick’(M) and f_3 links Travel(M) to Epidemic’.
Since an action can be applied to each person concurrently, the amount of possible actions is exponential due to the power set, i.e., all possible combinations for all persons. Foreplan avaoids the exponential blow-up by exploiting the indistinguishability of action PRVs.
With Definition 3, we require the rewards to be represented as a sum of local reward functions, and not as some arbitrary function. Actually, this is too restrictive: It is sufficient that an operation running over same behaving variables has a hyperoperation repeating the operation multiple times. For a sum, the hyperoperation is multiplication. Also, multiple terms of such operations can be linked together in any way. In Appendix A.1, we give examples for hyperoperations. For simplicity and the proofs, we use the definition as it is throughout the paper.
We exploit the symmetries in Definition 3 with Foreplan in the next section. In the remainder of this section, we explain what we mean by action PRVs and by parameterized local reward functions.